# Gradient as Conditions: Rethinking HOG for All-in-one Image Restoration

### v2

@@ Introduction
We evaluate HOGformer in two standard AIR settings: (I) 3-task adverse weather removal (Sun et al. 2024), including desnowing, draining&dehazing, and raindrop removal; and (II) 5-task general restoration (Zheng et al. 2024), including deraining, low-light enhancement, desnowing, dehazing, and deblurring. One AIR model is trained for each setting.
Datasets and Metrics. For Setting (I), we use the AllWeather dataset (Sun et al. 2024). For Setting (II), we use task-specific datasets: Merged rain dataset (Jiang et al. 2020; Wang et al. 2022b) for deraining, LOL (Wei et al. 2018) for low-light, Snow100K (Liu et al. 2018) for desnowing, RESIDE (Li et al. 2018a) for dehazing, and GoPro (Nah, Hyun Kim, and Mu Lee 2017) for deblurring. To evaluate generalization, we additionally test on real-world datasets: Practical (Yang et al. 2017), MEF (Ma, Zeng, and Wang 2015), NPE (Wang et al. 2013), DICM (Lee, Lee, and Kim 2013), HIDE (Shen et al. 2019), RealBlur (Rim et al. 2020), and T-OLED (Zhou et al. 2021).

### v3

@@ Introduction
Dehazing Deraining Denoising on BSD68 Average
Method on SOTS on Rain100L $\boldsymbol{\sigma=15}$ $\boldsymbol{\sigma=25}$ $\boldsymbol{\sigma=50}$
P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$
BRDNet 23.23 0.895 27.42 0.895 32.26 0.898 29.76 0.836 26.34 0.693 27.80 0.843
LPNet 20.84 0.828 24.88 0.784 26.47 0.778 24.77 0.748 21.26 0.552 23.64 0.738
FDGAN 24.71 0.929 29.89 0.933 30.25 0.910 28.81 0.868 26.43 0.776 28.02 0.883
MPRNet 25.28 0.955 33.57 0.954 33.54 0.927 30.89 0.880 27.56 0.779 30.17 0.899
DL 26.92 0.931 32.62 0.931 33.05 0.914 30.41 0.861 26.90 0.740 29.98 0.876
AirNet 27.94 0.962 34.90 0.968 33.92 0.933 31.26 0.888 28.00 0.797 31.20 0.910
PromptIR 30.58 0.974 36.37 0.972 33.98 0.933 31.31 0.888 28.06 0.799 32.06 0.913
AdaIR 31.06 0.980 38.64 0.983 34.12 0.935 31.45 0.892 28.19 0.802 32.69 0.918
HOGformer-L 31.91 0.981 38.50 0.983 34.04 0.935 31.40 0.892 28.16 0.804 32.80 0.919
Table 3: Comparisons under the three-degradation all-in-one setting.
Dehazing Deraining Denoising Deblurring Low-Light Average
Method on SOTS on Rain100L on BSD68 on GoPro on LOL
P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$ P $\uparrow$ S $\uparrow$
NAFNet 25.23 0.939 35.56 0.967 31.02 0.883 26.53 0.808 20.49 0.809 27.76 0.881
HINet 24.74 0.937 35.67 0.969 31.00 0.881 26.12 0.788 19.47 0.800 27.40 0.875
MPRNet 24.27 0.937 38.16 0.981 31.35 0.889 26.87 0.823 20.84 0.824 28.27 0.890
DGUNet 24.78 0.940 36.62 0.971 31.10 0.883 27.25 0.837 21.87 0.823 28.32 0.891
MIRNetV2 24.03 0.927 33.89 0.954 30.97 0.881 26.30 0.799 21.52 0.815 27.34 0.875
SwinIR 21.50 0.891 30.78 0.923 30.59 0.868 24.52 0.773 17.81 0.723 25.04 0.835
Restormer 24.09 0.927 34.81 0.962 31.49 0.884 27.22 0.829 20.41 0.806 27.60 0.881
DL 20.54 0.826 21.96 0.762 23.09 0.745 19.86 0.672 19.83 0.712 21.05 0.743
TransWeather 21.32 0.885 29.43 0.905 29.00 0.841 25.12 0.757 21.21 0.792 25.22 0.836
TAPE 22.16 0.861 29.67 0.904 30.18 0.855 24.47 0.763 18.97 0.621 25.09 0.801
AirNet 21.04 0.884 32.98 0.951 30.91 0.882 24.35 0.781 18.18 0.735 25.49 0.846
IDR 25.24 0.943 35.63 0.965 31.60 0.887 27.87 0.846 21.34 0.826 28.34 0.893
AdaIR 30.53 0.978 38.02 0.981 31.35 0.889 28.12 0.858 23.00 0.845 30.20 0.910
HOGformer-L 31.16 0.979 38.05 0.981 31.19 0.884 28.62 0.867 24.46 0.858 30.70 0.914
Table 4: Comparisons for five-degradation all-in-one restoration. Denoising results are reported at a noise level of $\sigma=25$. Methods in the top super-row represent general image restoration methods, while the remaining methods are all-in-one methods.
@@
We evaluate HOGformer in four standard AIR settings: (I) 3-task adverse weather removal (Sun et al. 2024), including desnowing, draining&dehazing, and raindrop removal; and (II) 5-task general restoration (Zheng et al. 2024), including deraining, low-light enhancement, desnowing, dehazing, and deblurring. (III) 3-task general restoration following (Potlapalli et al. 2023; Cui et al. 2025). (IV) 5-task general restoration following (Potlapalli et al. 2023; Cui et al. 2025). One AIR model is trained for each setting.
Datasets and Metrics. For Setting (I), we use the AllWeather dataset (Sun et al. 2024). For Setting (II), we use task-specific datasets: Merged rain dataset (Jiang et al. 2020; Wang et al. 2022b) for deraining, LOL (Wei et al. 2018) for low-light, Snow100K (Liu et al. 2018) for desnowing, RESIDE (Li et al. 2018a) for dehazing, and GoPro (Nah, Hyun Kim, and Mu Lee 2017) for deblurring. To evaluate generalization, we additionally test on real-world datasets: Practical (Yang et al. 2017), MEF (Ma, Zeng, and Wang 2015), NPE (Wang et al. 2013), DICM (Lee, Lee, and Kim 2013), HIDE (Shen et al. 2019), RealBlur (Rim et al. 2020), and T-OLED (Zhou et al. 2021). For Settings (III) and (IV), we adopt the same datasets as those used in prior studies (Potlapalli et al. 2023; Cui et al. 2025).
@@
Setting (III). Table 3 presents quantitative comparisons under the three-degradation all-in-one setting, including dehazing, deraining, and denoising with different noise levels. HOGformer achieves the best average PSNR and SSIM among all methods, and consistently delivers top or near-top performance across all degradation types and noise levels. Compared with recent all-in-one restoration methods such as PromptIR and AdaIR, HOGformer demonstrates superior overall restoration capability and robustness, highlighting its effectiveness in handling multiple degradations within a unified framework.
Setting (IV). Table 4 reports quantitative comparisons of general image restoration methods and all-in-one methods across five degradations, including dehazing, deraining, denoising, deblurring, and low-light enhancement. Among all-in-one approaches, HOGformer consistently achieves the best or highly competitive PSNR and SSIM on all datasets, and obtains the highest average performance overall. Compared with both general restoration models and recent AIR methods, HOGformer demonstrates superior restoration quality across diverse degradations, highlighting its strong generalization ability and balanced effectiveness.